ALABAMA CORPORATION

…. manufactures a single product, cleverly named Product X. The following information is available for the calendar year 2018 just completed, during which they produced and sold 200,000 units. Sales for the year was $2,400,000. During the year, the company paid a sales commission of 5 percent of sales. The corporate income tax rate was 30%. Also assume that the per-unit manufacturing cost in 2018 was the same as in the previous year, 2017.

Raw materials purchases

$300,000

Direct labor

140,000

Depreciation - factory equipment

45,000

Depreciation - factory building

30,000

Depreciation - headquarters building

50,000

Factory insurance

15,000

Property taxes:

Factory

20,000

Headquarters

18,000

Utilities - factory

34,000

Utilities - sales/other

1,800

Administrative salaries

150,000

Indirect labor salaries

156,000

Sales office salaries

90,000

Inventory Information

12/31/2017

12/31/2018

Raw materials

$124,000

$152,000

Work in process

124,000

130,000

Finished Goods

109,000

119,000

REQUIRED:

Post all of the above a summary entries in T accounts.

Compute the direct materials used in production.

Compute the prime cost.

Compute the conversion cost.

Prepare the Schedule of Raw Materials Used

Prepare the Schedule of Cost of Goods Manufactured

Prepare the Schedule of Cost of Goods Sold

Calculate the manufacturing cost per unit

Prepare an income statement for 2018, in good form.

As part of the annual audit you notice that “Indirect labor salaries” included the $20,000 salary paid to Ken, who actually works in the advertising department of the company.

Does this seem reasonable? Why or why not?

Does this make a difference in terms of the net operating income? Why or why not?

What incentive would management have to include the advertising costs as “indirect labor”? (That is, what would their motivation be to do this?)
Answer #1

As per policy, only four parts of a question is allowed to answer, but more than four parts are answered here :

t-ACCOUNTS:
Raw Material Account
OB 124000 WIP 272000
PURCH. 300000 CB 152000
Total 424000 Total 424000
WIP Account
OB 124000 Finished Goods 706000
DM 272000 CB 130000
DL 140000
F O/h 300000
Total 836000 Total 836000
Finished Goods Account
OB 109000 COGS 696000
WIP 706000 CB 119000
Total 815000 Total 815000
Income Summary:
Sales 2400000
Less:COGS 696000
GP 1704000
Less: Expenses:
Adminis 218000
Selling 211800
Total expenses 429800
IBT 1274200
Tax 382260
IAT 891940
Factory Overhead Account
Ind. Labor 156000 WIP 300000
Dep-factory 45000
Dep-fac Budg 30000
Insur 15000
Prop tax - fact 20000
Utilities 34000
Total 300000 Total 300000
Administration Exp Account
Dep - Budg 50000 Income Summary 218000
Prop tax-HQ 18000
Salaries 150000
Total 218000 Total 218000
Selling Exp Account
Utilities 1800 Income Summary 211800
salaries 90000
Commission 120000
Total 211800 Total 211800
REQ 1) RAW MATERIAL USED = $272000
REQ 2) PRIME COST = DM + DL = 272000+140000= 412000
REQ 3) CONVERSION COST = DL+factory overhead = 140000+300000=440000
REQ 4) SCHEDULE OF RAW MATERIAL USED:
OB 124000
add: PURCHASES 300000
Total available 424000
less: CB 152000
RM USED 272000
REQ 5) SCHEDULE OF COGM :
RM USED 272000
DL USED 140000
F OVERHEAD 300000
Total Cost incurred 712000
Add: WIP OB 124000
Total WIP 836000
Less : WIP CB 130000
COGM 706000
REQ 6) SCHEDULE OF COGS :
FG OB 109000
add: COGM 706000
Total FG available 815000
less: FG CB 119000
COGS 696000
REQ 7) Manufacturing Cost per unit:
COGM / UNITS MANUFACTURED
706000 / 200000 = $3.53 PER UNIT